# BrainOmni: A Brain Foundation Model for Unified EEG and MEG Signals

### v1

@@ Abstract
Electroencephalography (EEG) and magnetoencephalography (MEG) measure neural activity non-invasively by capturing electromagnetic fields generated by dendritic currents. Although rooted in the same biophysics, EEG and MEG exhibit distinct signal patterns, further complicated by variations in sensor configurations across modalities and recording devices. Existing approaches typically rely on separate, modality- and dataset-specific models, which limits the performance and cross-domain scalability. This paper proposes BrainOmni, the first brain foundation model that generalises across heterogeneous EEG and MEG recordings. To unify diverse data sources, we introduce BrainTokenizer, the first tokenizer that quantises spatiotemporal brain activity into discrete representations. Central to BrainTokenizer is a novel Sensor Encoder that encodes sensor properties such as spatial layout, orientation, and type, enabling compatibility across devices and modalities. Building upon the discrete representations, BrainOmni learns unified semantic embeddings of brain signals by self-supervised pretraining. To the best of our knowledge, it is the first foundation model to support both EEG and MEG signals, as well as the first to incorporate large-scale MEG pretraining. A total of 1,997 hours of EEG and 656 hours of MEG data are curated and standardised from publicly available sources for pretraining. Experiments show that BrainOmni outperforms both existing foundation models and state-of-the-art task-specific models on a range of downstream tasks. It also demonstrates strong generalisation to unseen EEG and MEG devices. Further analysis reveals that joint EEG-MEG (EMEG) training yields consistent improvements across both modalities. Code and model checkpoints will be released upon acceptance.
@@ 2.1 EMEG Signals
[…]
@@ 2.2 BrainTokenizer
The temporal representation $\mathbf{Z}_{\text{time}}$ and the sensor embedding $\mathbf{V}$ are then fed into a specialised cross-attention block, which performs channel compression to unify signals with varying numbers of channels into a fixed number of latent source variable. The Value of the cross-attention block is set to the temporal representation, Key is the sum of temporal representation and sensor embedding, and a set of learnable Query is used to adaptively aggregate information from temporal and spatial representations, producing intermediate latent features $\mathbf{Z}_{\text{src}}\in\mathbb{R}^{C^{\prime}\times W\times D}$, where $C^{\prime}$ is the number of latent source variables determined as a hyperparameter.
Quantization.
@@
$\mathcal{L}_{\text{time}}=||\mathbf{X}-\hat{\mathbf{X}}||,$ (1)
where |||||$\cdot$||| | ⋅ | | denotes L1 distance; (ii) a frequency-domain loss between the original and reconstructed amplitude spectrum $\mathbf{A}$ and phase spectrum $\mathbf{\Phi}$:
$\mathcal{L}_{\text{freq}}=||\mathbf{A}-\hat{\mathbf{A}}||+||\mathbf{\Phi}-\hat{\mathbf{\Phi}}||;$ (2)
@@ 3.1 Pretraining
1997 hours of EEG data and 656 hours of MEG data collected from open-source datasets are used for pretraining (see Appendix D for details), including EEG devices with channel numbers ranging from 19 to 128 and MEG devices with channel numbers ranging from 157 to 306. Among them, 85% of the data is used for training, 10% of the data is used for validation, and 5% of the data is used for testing. Additionally, one EEG dataset and one MEG dataset, which were both collected with unique device systems different from those in the training data, were excluded from the training data to evaluate the model’s cross-device generalisation ability.
@@
Minimal and standardised preprocessing was applied to maximise data utilisation. A 0.1Hz-96 Hz bandpass filter and 50/60 notch filters were first applied to remove slow drifts and power-line noise, and MEG recordings with HPI coils were additionally filtered in their HPI frequency bands. All signals were then resampled to 256 Hz. Bad channels were identified via a power-spectral-density-based detection algorithm and subsequently interpolated. To address reference inconsistencies and wide amplitude variations, reference values are first subtracted across all sensors of each type within each sample. Then, each channel is normalised to zero mean and unit variance at the sample level. Sensor coordinates and orientations are obtained from either the dataset-provided positions or the device’s standard montage.
@@ 3.2 Evaluation
Motor Imagery PhysioNet-MI Schalk2004 109 64 4.0 19571 4
@@
Ten different datasets for nine separate tasks are collected to evaluate the performance of BrainOmni. Overview of datasets, including modality, tasks, number of subjects, channels, segments, categories, and duration of each segment, are listed in Table 1, and details are listed in Appendix E. To prevent information leakage, all datasets used for downstream testing were excluded from the pretraining process. Consistent with pretraining, bandpass filtering, notch filtering, resampling and normalisation were applied for downstream dataset preprocessing.
@@
The proposed method is compared to four specialised models without pretraining on both EEG and MEG data: CNN-Transformer Wei2022 , ContraWR Yang2023contrawr SPaRCNet PMID:36878708 and ST-Transformer sttransformer . Two additional EEG foundation models are included for EEG data: LaBraM jiang2024large and CBraMod wang2025cbramod . Details of these baselines can be found in Appendix F.
To ensure fair comparison, BrainOmni is compared to all baselines using the same preprocessing procedures, training pipeline and downstream evaluation strategy. The only exception is that in the preprocessing of LaBraM and CBraMod, we followed the same setting of sampling rate and bandpass filter as their preprocessing. All experiments used a fixed random seed 42 and trained for 30 epochs to ensure reproducibility. Cross-subject 3-fold cross-validation was conducted, and the mean and standard deviation of three runs were reported. The model checkpoint that achieved the highest validation-set BACC was selected for testing. Additionally, experiments were conducted with learning rates of $[3\times 10^{-6},1\times 10^{-5},3\times 10^{-5}]$ respectively for each model and chose the learning rate yielding the best validation performance. Given class imbalances in some datasets, balanced accuracy (BACC) was chosen as the primary evaluation metric.
@@ 4.1 Downstream Evaluation
CNN-Transformer Wei2022 3.2M 0.641 $\pm$ 0.100 0.709 $\pm$ 0.148 0.513 $\pm$ 0.028 0.809 $\pm$ 0.004
ContraWR Yang2023contrawr 1.6M 0.609 $\pm$ 0.038 0.706 $\pm$ 0.067 0.591 $\pm$ 0.011 0.806 $\pm$ 0.005
SPaRCNet PMID:36878708 1.3M 0.563 $\pm$ 0.057 0.764 $\pm$ 0.048 0.496 $\pm$ 0.124 0.789 $\pm$ 0.002
ST-Transformer sttransformer 2.6M 0.663 $\pm$ 0.011 0.786 $\pm$ 0.001 0.535 $\pm$ 0.028 0.811 $\pm$ 0.008
LaBraM jiang2024large 5.8M 0.715 $\pm$ 0.065 0.843 $\pm$ 0.042 0.600 $\pm$ 0.109 0.807 $\pm$ 0.003
CBraMod wang2025cbramod 4.9M 0.649 $\pm$ 0.006 0.811 $\pm$ 0.006 0.617 $\pm$ 0.162 0.773 $\pm$ 0.001
BrainOmni_tiny 8.4M 0.800 $\pm$ 0.035 0.834 $\pm$ 0.036 0.658 $\pm$ 0.062 0.816 $\pm$ 0.001
BrainOmni_base 33M 0.806 $\pm$ 0.053 0.841 $\pm$ 0.044 0.733 $\pm$ 0.102 0.813 $\pm$ 0.007
@@
CNN-Transformer Wei2022 3.2M 0.312 $\pm$ 0.012 0.353 $\pm$ 0.008 0.618 $\pm$ 0.015 0.309 $\pm$ 0.009
ContraWR Yang2023contrawr 1.6M 0.342 $\pm$ 0.036 0.351 $\pm$ 0.007 0.509 $\pm$ 0.007 0.336 $\pm$ 0.033
SPaRCNet PMID:36878708 1.3M 0.357 $\pm$ 0.032 0.375 $\pm$ 0.010 0.738 $\pm$ 0.017 0.367 $\pm$ 0.015
ST-Transformer sttransformer 2.6M 0.374 $\pm$ 0.023 0.406 $\pm$ 0.008 0.782 $\pm$ 0.014 0.367 $\pm$ 0.035
LaBraM jiang2024large 5.8M 0.596 $\pm$ 0.016 0.444 $\pm$ 0.015 0.819 $\pm$ 0.024 0.372 $\pm$ 0.011
CBraMod wang2025cbramod 4.9M 0.416 $\pm$ 0.042 0.430 $\pm$ 0.021 0.532 $\pm$ 0.005 0.313 $\pm$ 0.009
BrainOmni_tiny 8.4M 0.597 $\pm$ 0.019 0.454 $\pm$ 0.012 0.818 $\pm$ 0.017 0.434 $\pm$ 0.006
BrainOmni_base 33M 0.618 $\pm$ 0.034 0.466 $\pm$ 0.008 0.819 $\pm$ 0.022 0.439 $\pm$ 0.011
Table 3: Baseline comparison on the ASD74 (MEG) and SomatoMotor (EMEG). Balanced accuracy (BACC) and F1 score are reported in mean $\pm$ standard deviation. “PT” stands for whether the model involves pretraining. The best results are shown in bold and second-best results are underlined.
# Param PT ASD74 SomatoMotor
BACC F1 BACC F1
CNN-Transformer Wei2022 3.2M 0.508 $\pm$ 0.011 0.069 $\pm$ 0.074 0.508 $\pm$ 0.010 0.524 $\pm$ 0.066
ContraWR Yang2023contrawr 1.6M 0.533 $\pm$ 0.047 0.343 $\pm$ 0.246 0.505 $\pm$ 0.003 0.233 $\pm$ 0.222
SPaRCNet PMID:36878708 1.3M 0.565 $\pm$ 0.039 0.463 $\pm$ 0.047 0.548 $\pm$ 0.008 0.543 $\pm$ 0.037
ST-Transformer sttransformer 2.6M 0.532 $\pm$ 0.024 0.298 $\pm$ 0.122 0.624 $\pm$ 0.063 0.646 $\pm$ 0.094
BrainOmni_tiny 8.4M 0.606 $\pm$ 0.019 0.526 $\pm$ 0.024 0.787 $\pm$ 0.112 0.794 $\pm$ 0.105
BrainOmni_base 33M 0.624 $\pm$ 0.024 0.561 $\pm$ 0.047 0.801 $\pm$ 0.128 0.788 $\pm$ 0.150
BrainOmni was evaluated on downstream EEG, MEG, and EMEG tasks. Results on EEG datasets are shown in Table 2. Results on MEG and EMEG datasets are listed in Table 3. It can be observed that BrainOmni achieved the highest performance on nearly all tasks, with a close second-best place on the MDD dataset. Specifically, on EEG data, BrainOmni_base’s balanced accuracy on the AD65 and PD31 datasets was about 10% higher than the best pretrained baseline LaBraM. On the MEG ASD74 dataset, it exceeded the strongest baseline by about 7%, and on the EMEG SomatoMotor task, it outperformed the top baseline by 20%. The results demonstrate that BrainOmni consistently outperforms both specialised and other pretrained foundation models across diverse downstream tasks and modalities, including EEG, MEG, and EMEG data, exhibiting strong generalisation and versatility.
@@ 4.2 Cross Device Generalisation
Table 4: Zero-shot reconstruction results of BrainTokenizer on unseen devices. “AMP” stands for amptitude MAE. PHASE stands for phase MAE. $\uparrow$” indicates the higher the better. $\downarrow$” indicates that the lower the better.
@@
To evaluate BrainTokenizer’s generalisation to entirely unseen EEG and MEG device systems, we selected two datasets, PerceiveImagine (EEG, SynAmps2 system) and Gloups-MEG (MEG, NeuroImaging system), whose recording setups were not included in any pretraining data. We compare the zero-shot reconstruction losses of these unseen datasets against those of the known-device EEG and MEG test set whose training partitions have been used in training BrainTokenizer.
Results are listed in Table 4. For EEG, the zero-shot results on PerceiveImagine even consistently outperform our EEG test set across all metrics, highlighting strong generalisation to a new EEG device system. For MEG, although the reconstruction losses of BrainTokenizer on Gloups-MEG are slightly worse than those of our own MEG test set, the relatively strong PCC value of 0.695 still indicates the model’s competitive generalisation capability for MEG devices. The slight performance degradation on MEG devices may be attributed to the relatively limited amount of MEG training data.
@@ 4.3 EMEG Joint Pretraining
EEG only 0.807 $\pm$ 0.006 0.555 $\pm$ 0.003 0.737 $\pm$ 0.042
MEG only 0.530 $\pm$ 0.032 0.683 $\pm$ 0.168
EMEG 0.816 $\pm$ 0.001 0.596 $\pm$ 0.019 0.606 $\pm$ 0.019 0.747 $\pm$ 0.088 0.786 $\pm$ 0.121 0.787 $\pm$ 0.112
@@
BrainOmni 0.816 $\pm$ 0.001 0.597 $\pm$ 0.019 0.800 $\pm$ 0.035 0.606 $\pm$ 0.019 0.787 $\pm$ 0.112
–w/o SE 0.785 $\pm$ 0.007 0.604 $\pm$ 0.018 0.730 $\pm$ 0.056 0.549 $\pm$ 0.007 0.643 $\pm$ 0.066
Pure temporal 0.786 $\pm$ 0.003 0.538 $\pm$ 0.010 0.774 $\pm$ 0.060 0.575 $\pm$ 0.031 0.740 $\pm$ 0.099
To investigate the effect of joint EEG and MEG training on model performance, two additional BrainOmni variants were trained using only EEG or only MEG data. The variants are compared to the BrainOmni model on EEG datasets TUAB, TUEV, MEG dataset ASD74, and the multimodal SomatoMotor dataset. Specifically for the SomatoMotor datasets, downstream experiments were conducted with EEG-only, MEG-only, and combined EMEG inputs separately. As shown in Table 5, joint EMEG pretraining consistently outperforms single-modality pretraining across all datasets. Notably, for the MEG modality of the SomatoMotor dataset, the jointly pretrained EMEG model improves balanced accuracy by 10% compared with the MEG-only model. The results highlight the benefit of multimodal training, particularly for the MEG modality with less accessible pretraining data. Additionally, comparing the EEG, MEG, and EMEG inputs processed by the EMEG model on SomatoMotor, the combined EMEG input achieves better performance than both EEG and MEG inputs, underscoring the effectiveness of jointly leveraging EEG and MEG data in downstream tasks.
(a) Learning curve
(b) Different codebook layers
(c) Latent source variables
Figure 3: (a) The training loss and validation loss curves during the training phase of BrainTokenizer. (b) The accuracy curves for parallel mask prediction on the labels of each codebook layer during the training phase of BrainOmni on the training set. (c) Trend of signal reconstruction loss over the number of latent source variables.

### v3

@@ Abstract
Electroencephalography (EEG) and magnetoencephalography (MEG) measure neural activity non-invasively by capturing electromagnetic fields generated by dendritic currents. Although rooted in the same biophysics, EEG and MEG exhibit distinct signal patterns, further complicated by variations in sensor configurations across modalities and recording devices. Existing approaches typically rely on separate, modality- and dataset-specific models, which limits the performance and cross-domain scalability. This paper proposes BrainOmni, the first brain foundation model that generalises across heterogeneous EEG and MEG recordings. To unify diverse data sources, we introduce BrainTokenizer, the first tokenizer that quantises spatiotemporal brain activity into discrete representations. Central to BrainTokenizer is a novel Sensor Encoder that encodes sensor properties such as spatial layout, orientation, and type, enabling compatibility across devices and modalities. Building upon the discrete representations, BrainOmni learns unified semantic embeddings of brain signals by self-supervised pretraining. To the best of our knowledge, it is the first foundation model to support both EEG and MEG signals, as well as the first to incorporate large-scale MEG pretraining. A total of 1,997 hours of EEG and 656 hours of MEG data are curated and standardised from publicly available sources for pretraining. Experiments show that BrainOmni outperforms both existing foundation models and state-of-the-art task-specific models on a range of downstream tasks. It also demonstrates strong generalisation to unseen EEG and MEG devices. Further analysis reveals that joint EEG-MEG (EMEG) training yields consistent improvements across both modalities. Code and checkpoints are publicly available at https://github.com/OpenTSLab/BrainOmni.
@@ 2.1 EMEG Signals
EEG measures electrical potentials through electrodes placed on the scalp, while MEG uses either gradiometer (GRAD) or magnetometer (MAG) sensors to measure the magnetic field at specific locations and orientations outside the head. The measured EMEG signals are multichannel time-series data, which can be represented as $\mathbf{X}\in\mathbb{R}^{C\times T}$, where $C$ denotes the number of sensor channels, and $T$ denotes the number of sampling points. Taking the sensor type, locations and orientations into account, an EMEG sample can be donated as $\boldsymbol{\Omega}=(\mathbf{X},\mathbf{L},\mathbf{S})$, where $\mathbf{X}$ is the raw signal, $\mathbf{L}\in\mathbb{R}^{C\times 6}$ is the physical position and orientation information of each sensor in Cartesian coordinates, and $\mathbf{S}=\{s_{1},s_{2},\dots,s_{c}|s_{i}\in\{0~\text{(EEG)},1~\text{(GRAD)},2~\text{(MAG)}\}\}$ is the type of each sensor.
@@ 2.2 BrainTokenizer
The temporal representation $\mathbf{Z}_{\text{time}}$ and the sensor embedding $\mathbf{V}$ are then fed into a specialised cross-attention block, which performs channel compression to unify signals with varying numbers of channels into a fixed number of latent source variable. The Value of the cross-attention block is set to the temporal representation, Key is the sum of temporal representation and sensor embedding, and a set of learnable Query is used to adaptively aggregate information from temporal and spatial representations, producing intermediate latent features $\mathbf{Z}_{\text{src}}\in\mathbb{R}^{C^{\prime}\times W\times D}$, where $C^{\prime}$ is the number of latent source variables determined as a hyperparameter. The cross attention operation is performed independently for each window, ensuring the consistency of channel compression between different temporal windows.
Quantisation.
@@
[…]
where $||\cdot||$ denotes L1 distance; (ii) a frequency-domain loss between the original and reconstructed amplitude spectrum $\mathbf{A}$ and phase spectrum $\mathbf{\Phi}$:
[…]
@@ 3.1 Pretraining
1997 hours of EEG data and 656 hours of MEG data collected from open-source datasets are used for pretraining (see Appendix H for details), including EEG devices with channel numbers ranging from 19 to 128 and MEG devices with channel numbers ranging from 157 to 306. Among them, 85% of the data is used for training, 10% of the data is used for validation, and 5% of the data is used for testing. Additionally, one EEG dataset and one MEG dataset, which were both collected with unique device systems different from those in the training data, were excluded from the training data to evaluate the model’s cross-device generalisation ability.
@@
Minimal and standardised preprocessing was applied to maximise data utilisation. A 0.1Hz-96 Hz bandpass filter and 50/60 notch filters were first applied to remove slow drifts and power-line noise, and MEG recordings with HPI coils were additionally filtered in their HPI frequency bands. All signals were then resampled to 256 Hz. Bad channels were identified via a power-spectral-density-based detection algorithm (see Appendix E for details) and subsequently interpolated. To address reference inconsistencies and wide amplitude variations, reference values are first subtracted across all sensors of each type within each sample. Then, each channel is normalised to zero mean and unit variance at the sample level. Sensor coordinates and orientations are obtained from either the dataset-provided positions or the device’s standard montage.
@@ 3.2 Evaluation
Motor Imagery PhysioNet-MI Schalk2004 109 64 4.0 9837 4
@@
Eleven different datasets for nine separate tasks are collected to evaluate the performance of BrainOmni. Overview of datasets, including modality, tasks, number of subjects, channels, segments, categories, and duration of each segment, are listed in Table 1, and details are listed in Appendix I. To prevent information leakage, all datasets used for downstream testing were excluded from the pretraining process. Consistent with pretraining, bandpass filtering, notch filtering, resampling and normalisation were applied for downstream dataset preprocessing.
@@
The proposed method is compared to four specialised EEG models: CNN-Transformer Wei2022 , ContraWR Yang2023contrawr SPaRCNet PMID:36878708 and ST-Transformer sttransformer , and a specialised MEG model: FAMED FAMED . The specialised models were implemented without pretraining. Two additional EEG foundation models are included for EEG data: LaBraM jiang2024large and CBraMod wang2025cbramod . Details of these baselines can be found in Appendix J.
To ensure fair comparison, BrainOmni is compared to all baselines using the same preprocessing procedures, training pipeline and downstream evaluation strategy. The only exception is that in the preprocessing of LaBraM and CBraMod, we followed the same setting of sampling rate and bandpass filter as their preprocessing. All experiments were conducted under a 5-fold cross-validation setup, where each split allocated three folds for training, one for validation, and one for testing. Each configuration was run under two random seeds. To evaluate the model’s generalisation across subjects, a strict cross-subject split strategy was applied to all datasets where subjects in the training set do not appear in the validation or test sets. Given class imbalances in some datasets, balanced accuracy (BACC) was chosen as the primary evaluation metric. The mean and standard deviation of all 10 runs were reported. Models were trained for 30 epochs, and the model checkpoint that achieved the highest validation-set BACC was selected for testing. More details of downstream training can be found in Appendix D.
@@ 4.1 Downstream Evaluation
CNN-Transformer Wei2022 - 0.695 $\pm$ 0.118 0.846 $\pm$ 0.096 0.536 $\pm$ 0.090 0.795 $\pm$ 0.024
ContraWR Yang2023contrawr - 0.660 $\pm$ 0.104 0.863 $\pm$ 0.068 0.521 $\pm$ 0.069 0.800 $\pm$ 0.012
SPaRCNet PMID:36878708 - 0.582 $\pm$ 0.085 0.809 $\pm$ 0.051 0.534 $\pm$ 0.081 0.779 $\pm$ 0.017
ST-Transformer sttransformer - 0.604 $\pm$ 0.080 0.835 $\pm$ 0.040 0.530 $\pm$ 0.081 0.793 $\pm$ 0.008
LaBraM jiang2024large 5.8M 0.711 $\pm$ 0.060 0.880 $\pm$ 0.037 0.659 $\pm$ 0.188 0.816 $\pm$ 0.006
CBraMod wang2025cbramod 4.9M 0.681 $\pm$ 0.040 0.871 $\pm$ 0.048 0.584 $\pm$ 0.127 0.808 $\pm$ 0.007
BrainOmni_tiny 8.4M 0.795 $\pm$ 0.030 0.886 $\pm$ 0.043 0.736 $\pm$ 0.116 0.819 $\pm$ 0.004
BrainOmni_base 33M 0.828 $\pm$ 0.030 0.877 $\pm$ 0.052 0.748 $\pm$ 0.139 0.819 $\pm$ 0.005
@@
CNN-Transformer Wei2022 - 0.331 $\pm$ 0.047 0.357 $\pm$ 0.011 0.666 $\pm$ 0.008 0.337 $\pm$ 0.019
ContraWR Yang2023contrawr - 0.371 $\pm$ 0.030 0.347 $\pm$ 0.011 0.579 $\pm$ 0.031 0.385 $\pm$ 0.026
SPaRCNet PMID:36878708 - 0.362 $\pm$ 0.036 0.377 $\pm$ 0.012 0.748 $\pm$ 0.018 0.456 $\pm$ 0.016
ST-Transformer sttransformer - 0.392 $\pm$ 0.032 0.401 $\pm$ 0.008 0.749 $\pm$ 0.007 0.398 $\pm$ 0.015
LaBraM jiang2024large 5.8M 0.588 $\pm$ 0.017 0.458 $\pm$ 0.014 0.831 $\pm$ 0.015 0.561 $\pm$ 0.015
CBraMod wang2025cbramod 4.9M 0.525 $\pm$ 0.021 0.441 $\pm$ 0.014 0.822 $\pm$ 0.016 0.595 $\pm$ 0.015
BrainOmni_tiny 8.4M 0.603 $\pm$ 0.024 0.472 $\pm$ 0.018 0.825 $\pm$ 0.015 0.580 $\pm$ 0.019
BrainOmni_base 33M 0.622 $\pm$ 0.028 0.490 $\pm$ 0.013 0.832 $\pm$ 0.013 0.590 $\pm$ 0.022
Table 3: Baseline comparison on the ASD74 (MEG), MEG-MMI (MEG) and SomatoMotor (EMEG). Balanced accuracy are reported in mean $\pm$ standard deviation. “PT” stands for whether the model involves pretraining. The best results are shown in bold and second-best results are underlined.
# Param PT ASD74 MEG-MMI SomatoMotor
FAMED FAMED - 0.555 $\pm$ 0.038 0.527 $\pm$ 0.100 0.500 $\pm$ 0.016
CNN-Transformer Wei2022 - 0.517 $\pm$ 0.034 0.549 $\pm$ 0.074 0.503 $\pm$ 0.014
ContraWR Yang2023contrawr - 0.502 $\pm$ 0.009 0.528 $\pm$ 0.035 0.491 $\pm$ 0.020
SPaRCNet PMID:36878708 - 0.604 $\pm$ 0.053 0.581 $\pm$ 0.109 0.541 $\pm$ 0.027
ST-Transformer sttransformer - 0.583 $\pm$ 0.034 0.561 $\pm$ 0.054 0.661 $\pm$ 0.047
BrainOmni_tiny 8.4M 0.621 $\pm$ 0.048 0.610 $\pm$ 0.057 0.863 $\pm$ 0.128
BrainOmni_base 33M 0.651 $\pm$ 0.043 0.604 $\pm$ 0.061 0.832 $\pm$ 0.064
BrainOmni was evaluated on downstream EEG, MEG, and EMEG tasks. Results on EEG datasets are shown in Table 2. Results on MEG and EMEG datasets are listed in Table 3. It can be observed that BrainOmni achieves the highest performance on all tasks, with a close second-best place on the PhysioNet-MI dataset. Specifically, on EEG data, BrainOmni_base’s balanced accuracy on the AD65 and PD31 datasets is about 10% higher than the best pretrained baseline LaBraM. On the MEG ASD74 dataset, it exceeds the strongest baseline by about 5%, and on the EMEG SomatoMotor task, it outperforms the top baseline by 20%. The results demonstrate that BrainOmni consistently outperforms both specialised and other pretrained foundation models across diverse downstream tasks and modalities, including EEG, MEG, and EMEG data, exhibiting strong generalisation and versatility.
@@ 4.2 Cross Device Generalisation
Table 4: Zero-shot reconstruction results of BrainTokenizer on unseen devices. “AMP” stands for amptitude MAE. PHASE stands for phase MAE. $\uparrow$” indicates the higher the better. $\downarrow$” indicates the lower the better.
@@
To evaluate BrainTokenizer’s generalisation to entirely unseen EEG and MEG device systems, we selected two datasets, PerceiveImagine ds005620:1.0.0 (EEG, SynAmps2 system) and Gloups-MEG ds005261:3.0.0 (MEG, NeuroImaging system), whose recording setups were not included in pretraining data. We compare the zero-shot reconstruction losses of these unseen datasets against those of the known-device EEG and MEG test set whose training partitions have been used in training BrainTokenizer.
Results are shown in Table 4. For EEG, the zero-shot results on PerceiveImagine even consistently outperform our EEG test set across all metrics, highlighting strong generalisation to a new EEG device system. For MEG, although the reconstruction losses of BrainTokenizer on Gloups-MEG are slightly worse than those of our own MEG test set, the relatively strong PCC value of 0.695 still indicates the model’s competitive generalisation capability for MEG devices. The slight performance degradation on MEG devices may be attributed to the relatively limited amount of MEG training data.
@@ 4.3 EMEG Joint Pretraining
EEG only 0.811 $\pm$ 0.006 0.583 $\pm$ 0.020 0.766 $\pm$ 0.070
MEG only 0.553 $\pm$ 0.062 0.802 $\pm$ 0.141
EMEG 0.819 $\pm$ 0.004 0.603 $\pm$ 0.024 0.621 $\pm$ 0.048 0.783 $\pm$ 0.076 0.838 $\pm$ 0.128 0.863 $\pm$ 0.128
@@
BrainOmni 0.819 $\pm$ 0.004 0.603 $\pm$ 0.024 0.795 $\pm$ 0.030 0.621 $\pm$ 0.048 0.863 $\pm$ 0.128
–w/o SE 0.784 $\pm$ 0.005 0.607 $\pm$ 0.016 0.730 $\pm$ 0.064 0.568 $\pm$ 0.070 0.744 $\pm$ 0.076
Pure temporal 0.788 $\pm$ 0.003 0.526 $\pm$ 0.016 0.763 $\pm$ 0.047 0.571 $\pm$ 0.053 0.766 $\pm$ 0.050
To investigate the effect of joint EEG and MEG training on model performance, two BrainOmni variants were trained using only EEG or only MEG data. The variants are compared to the BrainOmni model on EEG datasets (TUAB, TUEV), MEG dataset (ASD74), and multimodal dataset (SomatoMotor). For the SomatoMotor dataset, downstream experiments were conducted with EEG-only, MEG-only, and combined EMEG inputs separately. As shown in Table 5, joint EMEG pretraining consistently outperforms single-modality pretraining across all datasets. Notably, for the MEG ASD74 dataset, the jointly pretrained EMEG model shows a 12% relative BACC improvement compared with the MEG-only model. The results highlight the benefit of multimodal training, particularly for the MEG modality with less accessible pretraining data. Additionally, comparing the EEG, MEG, and EMEG inputs processed by the EMEG model on SomatoMotor, the combined EMEG input achieves better performance than both EEG and MEG inputs, underscoring the effectiveness of jointly leveraging EEG and MEG data in downstream tasks.
5 Analysis